### 05 Dec 2023 v1

# Chemical strategies to mitigate electrostatic charging during coffee grinding

Abstract
The process of grinding coffee generates particles with high levels of electrostatic charge, causing a number of detrimental effects such as clumping, particle dispersal, and spark discharges. At the brewing level, electrostatic aggregation between particles affects liquid-solid accessibility, leading to variable extraction quality. In this study, we quantify the effectiveness of four charge mitigation strategies. Our data suggests that adding small amounts of water to whole beans pre-grinding, or bombarding the grounds with ions produced from a high-voltage ionizer, are capable of de-electrifying the granular flows. While these techniques helped reduce visible mess, only the static reduction through water inclusion was found to impact the brewing parameters in espresso format coffee. There, wetting coffee with than 0.05 mL / g resulted in a marked shift in particle size distribution, in part due to preventing clump formation and also liberating fine particles from sticking to the grinder. With all other variables kept constant, this shift resulted at least 15% higher coffee concentration for espresso prepared with darker roasts. These findings have significant financial and sustainability implications, and encourage the widespread implementation of water use to de-electrify coffee during grinding.
keywords:
Coffee, Granular electrification, Static eliminationPACS:
0000, 1111MSC:
0000, 1111[inst1]organization=Electrical and Computer Engineering, Portland State University,addressline=1400 SW 4th Avenue, city=Portland, postcode=97201, state=Oregon, country=US; joshua.mendez@pdx.edu
[inst2]organization=Department of Chemistry and Biochemistry, University of Oregon, addressline=1253 University of Oregon, city=Eugene, postcode=97403, state=Oregon, country=US; chendon@uoregon.edu
1 Introduction
Grinding roasted coffee reduces whole beans into flows of highly electrified powders. lim_triboelectric_2021 ; mendez2023moisture Charged granular materials can lead to electrostatic discharges, eckhoff2007electrostatic jamming and sheeting (i.e. coating the interior walls of conduits), eilbeck2000effect ; hendrickson2006electrostatics spontaneous segregation,mehrotra2007spontaneous and product non-uniformity.elajnaf2006electrostatic Specific to coffee preparation, charging can produce erratic dispersal of grounds, making whole bean grinders somewhat messy. More importantly, however, static charging during grinding results in particle-particle clumping.lim_triboelectric_2021 ; mendez2023moisture These electrostatic agglomerates affect extraction quality when brewing, by changing the packing of coffee particles and influencing solid surface area available to percolating water.cameron_systematically_2020 ; gagne2020physics The elimination of these clumps poses to increase soluble availability by upwards of 15%, posing substantial financial and sustainability motivations to eliminate their formation.
Recently, we demonstrated parameters that control charging by grinding commercially-sourced coffees and measuring the charge-to-mass ratio of the grounds using the process presented in Figure 1a.mendez2023moisture As whole beans are fractured into small grains with broad size distributions (Figure 1b), uman2016effect particles may acquire charge densities comparable to those of volcanic ash and thundercloud ice, through both fracto- and triboelectrification.miura2002measurements ; mendez2021charge Generally, the polarity and magnitude of charge loosely depends on the roast level or color of a coffee, and we observed that dark roasts charge largely negative, and lighter roasts charging positively, Figure 1c. Residual moisture levels, a property typically inversely proportional to color, was found to be the primary determiner of charge polarity, where positive-to-negative charging occurs at moisture contents less than 2%. Some coffees, especially dark coffees, can charge sufficiently to cause gaseous breakdown in the form of spark discharges, and when ground into a metal receptacle can cause spark discharges the span millimeters, Figure 1d.
The coffee industry has long maintained an intuitive understanding that water can significantly modulate grinding-associated electrostatics. A small amount of water may be added to whole beans before grinding is known to prevent static accumulation, and causes the grinder to retain less grounds within its chamber.homebarista2012static However, the inclusion of too much water may result in caking or corrosion in the grinder, the limit of which has not been determined. Concurrently, there is interest in developing water-free charge mitigation strategies, but their utility and impact on acheiving high extraction yields remains unknown. In this paper, we examine the effectiveness of various electrochemical techniques to suppress electrostatic build up – grinding and waiting, adding external water, and using two chemical ionization methods. We show that de-electrification techniques that introduce charges after the agglomerates have formed (i.e. simply waiting or deionizing radiative methods) does not improve the particle availability in espresso brewing, resulting in variable extractions and no improvement in total dissolved solids (%TDS). The addition of water mitigates charging during the grinding process, and results in extraction increase of beyond 15%.
2 Methods
We electrified coffee by grinding whole beans using a Mahlkönig EK43 grinder with stock 98mm burrs. Building on our previous work that showed that internal water content, not origin or processing method, determines charging of roasted whole beans. Here we conducted experiments using three coffees (two Mexican coffees and one Ethiopian) roasted in-house on an Ikawa Pro100 roaster under two temperature profiles (see Figure S1), yielding dark and light colors. Salient characteristics of both green and roasted coffee are noted in Table 1. The coffee color/roast degree and internal water content were measured using The Dipper KN-201 and the RoastRite RM-800, respectively.
We assessed the performance of four techniques to reduce static generated during grinding: time-resolved discharge, the addition of external water, corona discharge, and radioactive ionization. With the exception of the grind-and-wait technique, all experiments consisted in applying an electrostatic reduction technique during grinding and measuring the residual charges on particles exiting the grinder using a Faraday cup. Charge-to-mass (Q/m) ratios were then calculated, allowing comparative analysis between the various techniques. The particulars of each method are described in the sections that follow.
Prior to grinding, roasted coffee was stored in sealed, evacuated bags and kept at 253 K. Prior to grinding the coffee was allowed to reach equilibrium temperature before unsealing. For all experiments, we used a grind setting of 2 (arbitrary units) on the EK43, producing particle size distributions comparable to those shown in Figure 1b. Particle size measurements were performed on a Malvern Mastersizer 2000 with the solid-particle feed system, Scirocco 2000. Given the wide conditions under which coffee is prepared, we conducted experiments in a room at 20C and 35% RH. Each Faraday cup experiment was conducted a minimum of 3 times.
| Yirgacheffe Zero Defect | Temascaltepec | Yogondoy | |
| Origin |
Ethiopia |
México |
México |
| Producer |
Tamrat
Federico
García Luna |
| Process |
Washed |
Washed |
Washed |
%H2O (initial) |
12.0 |
8.9 |
9.3 |
%H2O (dark) |
1.0 |
1.3 |
1.1 |
%H2O (light) |
2.8 |
3.0 |
3.0 |
Agtron (dark) |
62.1 |
58.4 |
60.2 |
Agtron (light) |
88.7 |
70.1 |
93.1 |
3 Canvasing charge passivation strategies
Time-resolved discharge
Perhaps the simplest discharge method is to let the ground coffee sit for a period of time after grinding. This respite permits discharge through volumetric or surface conduction or, in the case of exceedingly high charges, gaseous breakdown.paiva2019conduction ; mendez2022lifetime ; navarro2023surface To a zeroth order, Jones and Tang jones1987charge have described the relaxation of the volumetric charge density in a powder by;
where is the initial charge density, is time, is the dielectric constant of the material, is the permittivity of free space, and is the effective resistivity. The denominator in the exponent defines a time constant, . This exponential behavior of charge decay can be readily observed using a non-contact electrostatic voltmeter probe (Trek 541A-2) placed 5 mm over 10 g of freshly ground coffee (collected in a metal cup, resting on an insulating surface), Figure 2a. Charge relaxation curves for both light and dark roasts of the same Ethiopian coffee (a washed Yirgacheffe) are rendered in Figure 2b. There, the light roast (2.8 % residual water, higher resistivity) loses its charge faster ( 15 s) than its darker, drier counterpart (1.0 % residual water, 65 s). Overall, however, charge appears to dissipate on timescales of minutes, but exceed the average time between preparing shots in a busy cafe. Also grinding and waiting poses problems for volatile loss and quality degradation, which occurs over similar time frames.degas
De-electrification through external water inclusion
While time-resolved de-electrification is both cost-effective and predictable, it does not prevent the formation of aggregates during grinding and also provides the particles with prolonged off-gassing time, resulting in a significant loss of volatiles.smrke2017time Thus, a number of active strategies have been devised to address charging during the act of grinding. We previously demonstrated that extrinsic water mitigated fractocharging in both positive and negative charging coffees.mendez2023moisture By incorporating up to 30 L g-1 of whole coffee beans, our in-house roasts behave similarly to other literature coffee samples, with the charge-to-mass ratio decreasing with increasing water content, Figure 2c. Although, in many cases, charging was not completely eradicated, we find that 20-30 L g-1 reduced the charge by a minimum of 50-60 %. In practice, this reduction appears to sufficiently mute electrostatic forces, precluding dispersal of grains, clumping, and other effects. A video of the behaviour is presented in the Supplementary Information (Video S2mendez2023fig ).
Although we have not observed it in our hands, the addition of water could lead to residual water accumulation within the grinder. This could pose problems for bacterial growth within the chamber or corrosion of the burrs over time. By placing a small humidity sensor within the grinding volume of the EK43, we measured the build up of moisture associated with the water droplet technique. For additive water in the range of 0-50 l g-1 and a base RH of 40%, we find that the relative humidity (%RH) within the grinder may increase up to 75% for a few seconds (see Figure S2), but returns back to ambient within minutes. The water is presumably consumed in electrochemical reactions, or boiled off. We did not detect condensation. The grinding a 10-20 dry beans returns the grinder to equilibrium %RH instantaneously. We will return to the use of water as a charge mitigator later in the paper.
Static reduction using ion beams
To move away from adding water, charge may also be neutralized by recombination with extrinsically generated negative and positive ions. Such techniques draw from extensive heritage in other settings,revel2003generation ; pingali2009use ; wang2023NASA and generally employ one of two methods; 1) corona discharge, which uses high voltages at sharp, conducting tips to accelerate naturally-occurring free ions and also cause collisions with neutrals;kodama2002static ; steinman2006basics ; pingali2009use and 2) alpha ionization, involving a radioactive source emitting helium nuclei that then collide with neutral species, to generate similar numbers of positive and negative ions.czajkowski1999evaluation
Corona discharge – The efficacy of corona discharge on charge reduction can be tested using a high-voltage ionizer. The device consists of a sharp carbon tip placed some distance (30–120 mm) from the chute where the coffee exits the grinder (see schematic in Figure 3b). The tip is fed by a high-voltage source that can generate 12 kV. Using a Gerdien condenser (AIC, AlphaLabs Inc.), we estimate the negative and positive ion densities to be 1.5 and 1.2 106 m at a distance of 0.3 m, respectively. These densities can be augmented by moving the tip closer to the chute. To shield the Faraday cup from direct influence of generated ions, we placed a coarse, grounded copper mesh over the FC aperture.
The plots in Figure 3a show the Q/m ratios gained by dark and light roasts of the YirgZ and Yogondoy during grinding as a function of ion density (as measured by the Faraday cup). With the ionizer off, the dark roasts nominally charge negatively, whereas the light roasts gain positive charge. Increasing the positive ion density reduces the negative Q/m of the dark roasts toward 0 nC g (middle panel of Figure 3b). Any further increase causes a polarity flip and the dark roasts end up depositing positive charge into the Faraday cup. A similar effect if true for light roasts (rightmost panel of Figure 3b); low degrees of negative ionization reduce positive charge, but moderate-to-high ion densities result in negative charging.
Corona discharge is highly effective at neutralizing charge and minimizing dispersal effects, (i.e. mess, see Video S3), but only if the characteristics of the ion source are tuned to the charging behavior of a particular coffee. Despite originating from the same green coffee, the light and dark Yogondoy samples require vastly different ion densities and polarity to achieve a reduction in Q/m ratio comparable to that produced by the water addition technique. The dark roast necessitates ion densities around 6 - 7.5 106 m, whereas charge on the light roast is minimized between -5 and -2.5 106 m. While these ranges can be achieved by adjusting the distance between coffee and ion source (by moving the source 10-20 mm), ion densities outside of these ranges compound existing problems (and can even increase charge) and create new problems (such as scattering fine particles via an ionic wind). Because the behavior will depend on evironmental variables such as humidity, coffee moisture, roast color, and other parameters, implementing ionization discharge effectively would necessitates trial and error.
Radioactive static eliminators – For completeness, we also ground coffee in the presence of two commercially-available, radioactive static eliminators (RSE), Figure 3c. These RSEs (Staticmaster 2U500) make use of small amounts of 210Po to generate alpha particles with energies of 5.40 MeV. Widely used in powder technology applications, static eliminators are manufactured by encapsulating thin layers of bismuth in gold. Then, some of the bismuth is transmuted into polonium by bombarding the capsule with deuterons.younes2020production For these experiments, two facing RSEs were mounted at a distance of 20 mm from the chute as shown in Figure 3c. At the manufacture date, the RSEs had a combined, initial activity of 18.5 MBq (1 mCi). As the 210Po decays, emitted alpha particles generate both positive and negative ions, and electrons, in the vicinity of the eliminators. RSEs are balanced ionizers, meaning they generate roughly equal number of positive and negative charge carriers. In the space between the RSEs, we measured positive and negative ion densities in the range of 8-10 106 m (for a net density of 0 m-3).
During an experiment, ground coffee was allowed to fall through the volume of ionized air generated by the RSEs. The bar graphs in Figure 3d show the Q/m ratios measured on dark and light roasts of all three coffees in the absence (grey bars) and presence (green/yellow bars) of the static eliminators. For dark roasts (left panel), we observe a minimum reduction in charge-to-mass ratio of 50%, comparable to that imparted by the water droplet technique (at the highest water contents of 20-30 L g-1). For light roasts, we also see significant reductions in electrification. Interestingly, lightly-roasted samples that nominally charge positively (YirgZ and Yogondoy) may occasionally acquire small negative charges in the presence of the RSEs. We suspect this polarity flip reflects the higher mobility of electrons relative to that of much more massive positive ions. Note also that the bipolar RSE is overall less efficient than a properly tuned unipolar corona ionizer in its capacity to reduce charge (despite the fact that the RSE generates higher negative and positive ion densities). This observation is aligned with experiments in silos which show that the most effective neutralization of negatively-charged powders occurs not under balanced ionization, but when the positive ion density is higher than the negative one by a factor of 2-3.zhou2022electrostatic
While RSEs do not require power, their main disadvantage is the use of radioactive material. Beyond obvious health effects if improperly handled, 210Po has a half-life of 138 days, giving commercial RSEs working lives of around a year before most 210Po has decayed to stable 206Pb. Thus, while they effectively mitigate charging, RSEs have relatively short lifespans and we certainly are not advocating cafes source 210Po.
4 Aggregate formation
For the ionizing strategies, charge mitigation occurs at the chute exit, not in the grinder itself. Conversely, the water addition addresses charging throughout the process. These differences may have important effects on sheeting (i.e. wall adhesion), material loss, and clump formation, especially since previous work has shown that particles adhered to the wall can have Q/m ratios several dozen times larger than those forming the bulk.zhang2021experimental
Whether a particle of radius with charge will electrostatically adhere to a surface depends in part on its electrostatic-to-gravitational force ratio,
where is Coulomb’s constant, is the particle’s density, and is the acceleration due to gravity. Because , smaller particles are much more likely to adhere to surfaces than larger ones (for a given ). Thus, dry, charged coffee exiting the grinder tends to be depleted in small particles. Retained fines within the grinder must be knocked out mechanically. This segregation can be readily observed in Figure 4a, where we plot the size distribution of grounds directly expelled by the grinder (solid, brown curve) and that of grounds retained within the grinding cavity (dotted, brown curve) when 10 g of whole coffee is ground with no charge mitigation technique.
Our data shows that adding moisture to whole coffee reduces charging by at least 50%. This reduction would result in a 4-fold decrease in the , possibly allowing small particles to overcome electrostatic adhesion and become reincorporated in to the bulk. The effect is reflected in our data where the addition of 100 L of water added to 10 g of whole beans produces an appreciable shift in particle size distribution toward smaller diameters, Figure 4a (dashed, blue curves). In fact, we observe a decrease in the mean particle size of expelled grounds up to water contents of 50 l g-1 (see Figure 4b). However, at higher extrinsic moisture levels, the mean particle size again moves toward larger diameters due to the formation of moisture-promoted aggregation driven by capillary forces, rather than electrostatic ones.telling2013ash
For the high-voltage ionization system, we do not see an analogous shift toward smaller particles sizes with increasing ion density (see Figure 4c). Comparing the masses of coffee retained under high-voltage ionization and the water addition technique, Figure 4d, even a modest 10 l g-1 reduces retention to 2.5%, whereas the high-voltage ionization method has retention percentages indistinguishable from those of no-treatment grinding (12%).
5 The effect of charge mitigation on espresso quality
In principle, charge mitigation during grinding should provide a better control on the characteristics of the grounds are used during coffee brewing. However, as evident from Figure 4, disparate charge reduction methods, while generally effective at reducing a material’s Q/m ratio, do not necessarily generate granular materials with equatable properties. How do these differences influence brewing?
In our previous work,mendez2023moisture we demonstrated that adding small amounts of water to beans prior to grinding changed the brewing behaviors when preparing espresso. Figure 5a exemplifies such behaviors. There, we plot the shot time (left panel) and flow rate for espresso brewed with and without added water (all other espresso parameters were maintained the same: 18.0 g of dry mass coffee were used to produce 45.0 g of liquid coffee extract, ground at setting 1.0, tamped at 196 N, and brewed using 94 C water, kept at 7 bar of static water pressure with a 2 second pre-infusion on a Victoria Arduino Black Eagle). For these experiments, we employed the dark roasted Temascaltepec. Mitigating charge with extraneous water results in extended shot times, decreased flow rates, and increased percentages of total dissolved solids (%TDS) as compared to no charge mitigation. We interpreted these findings to reflect the breaking or rearrangement of electrostatically-bound aggregates, resulting in a particle bed with smaller average grain size and, thus, reduced permeability. The data in Figures 4a and d suggest that an additional mechanism may also be operative: added water dislodges a large number of small particles that would otherwise remain trapped within the grinder.
The impact of shifting the distribution finer is enormous. Keeping all variables the same, a remarkable 16% increase in coffee concentration is achieved. Such increase in accessible coffee material in comparable brew times, Figure 5a, poses significant financial implications for the coffee industry, allowing for more efficient use of dry mass coffee, at the cost of adding less than 0.5 mL of water to the whole beans during grinding.
Conversely, charge reduction using an ionizing source at the grinder chute does not produce a reduction in the mean particle size, nor does it decrease grinder retention (see Figures 4c and d). Thus, it is worth considering whether ionizer-based techniques produce departures from the nominal brewing behavior similar to those observed with added water. As evident in Figure 5b, we observe nearly indistinguishable differences in espresso shots prepared with or without the ionizers. The modest increase in %TDS from 7.76% to 8.02% (average over five replicates) is still notable, but the impact is far less pronounced. These data suggests that ionization techniques perhaps do cause some aggregates to break up as these exit the chute, but the effect is limited. However, they do not propitiate the reincorporation of small particles bound electrostatically to the interior surfaces of the grinder and burrs. Thus, unlike the water drop technique, discharging coffee at the nozzle produces limited changes to the physical characteristics of the grounds used to brew coffee compared to the untreated samples.
For dark coffees, added water prior to grinding can generate appreciable differences in espresso brew characteristics, Figure 5. Similar experiments with a lightly roasted sample of the same coffee (Temascaltepec) do not reveal significant changes to espresso brew characteristics, either with extrinsic water or ionization (Figure S3). As noted in our previous work,mendez2023moisture darker roasted coffees not only charge negatively, but also acquire the highest absolute charge-to-mass ratios. Lighter roasts charge more ineffectively, with coffees with 2% residual water acquiring Q/m ratios near 0 nC g-1. Consequently, electrostatic effects like clumping and sheeting are markedly less present when grinding lighter roasts. That the addition of extrinsic water or ionization do little to modify the properties of the bed is, thus, not surprising.
6 Conclusion
We have comprehensively assessed the performance of a broad range of electrostatic reduction techniques in the context of coffee grinding. Added moisture and high-voltage ionization effectively counteract charging generated through fracto- and triboelectric charging. At minimum, both techniques have the potential to decrease the gravimetric charge by at least 50 %. While ionization methods can reduce the Q/m ratios of expelled coffee to near zero if appropriately tuned to a given coffee, they generally do not address electrification processes and adhesion dynamics within the grinder. As such, ionization methods at the grinder chute do not mitigate material loss (retention) which, for a grinder like the EK43, can exceed 10%. Conversely, the charge reduction afforded by the water addition (involving volumes ranging from 0-50 L g-1 dry mass coffee) has the capacity to resolve aggregation effects across the entire grinding system, even if particles retain some charge. Because particle-wall aggregates comprise smaller grains, we find that reincorporating this material into the bulk significantly changes the espresso brewing behavior (slower shots and smaller flow rates) and resultant increase in coffee solubility. And while our experiments conclusively demonstrate the economic upside of water incorporation during grinding for espresso, we suspect a similar effect will be observed for all percolation brewing techniques, as liberation of fine particles will not only clog void space in the coffee bed, but also fill voids in filter paper in pour-over embodiments.
Acknowledgements
JMH acknowledges support from DOE Grant No. 242649. CHH acknowledges the support from the National Science Foundation under Grant No. 2237345 and support from the Camille and Henry Dreyfus Foundation. This work was supported by the Coffee Science Foundation, underwritten by Nuova Simonelli. The work was enabled by donations of green coffee from Farmers Union and Finca La Ilusión. We are grateful to Ikawa for providing the roaster, Pentair for providing the reverse osmosis water filtration system, and Tailored Coffee, Eugene, OR for the use of the EK43.